Earnings per Share, Price-Earnings Ratio, Dividend Yield

Earnings per Share, Price-Earnings Ratio, Dividend Yield The following information was taken from the financial statements of Tolbert Inc. for December 31 of the current fiscal year: Common stock, $10 par value (no change during the year) Preferred $10 stock, $100 par (no change during the year) The net income was $1,206,000 and the declared dividends on the common stock were $77,500 for the current year. The market price of the common stock is $20.80 per share. For the common stock, determine (a) the earnings per share, (b) the price-earnings ratio, (c) the dividends per share, and (d) the dividend yield. If required, round your answers to two $3,100,000 4,000,000 decimal places. a. Earnings per Share b. Price-Earnings Ratio c. Dividends per Share d. Dividend Yield
Answer #1

a) Earning per share = (1206000-400000)/310000 = $2.60 per share

b) Price earning ratio = MPS/EPS = 20.80/2.60 = 8 Times

c) Dividend per share = 77500/310000 = $0.25 per share

d) Dividend yield = 0.25*100/20.80 = 1.20%